OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch apothem A = sq. in. Please explain your answer

OpenStudy (sriram):

what exactly is the apothem?

OpenStudy (anonymous):

the perpendicular distance from the center of the polygon to a side.

OpenStudy (anonymous):

Isn't this just the same question as last time?

OpenStudy (anonymous):

No it has "apothem" now

OpenStudy (anonymous):

OK, not exactly...

OpenStudy (anonymous):

You can just use cos or sin to turn it into the same type of problem...

OpenStudy (sriram):

okay fine then the apothem will also be a part of the median & in properties of triangle it divides in the ratio 2:1 hence height is 18inches the side or side=18*tan30 hecce area=18*18*tan30* 1/2 =162*(1/3)^1/2

OpenStudy (sriram):

sorry i typed side or side it is side or base

OpenStudy (anonymous):

You see what I am saying, you have Sin 30 = apoth/circumradius. Once you have circumradius use the same formula as last time.

OpenStudy (anonymous):

How would I find the circumradius?

OpenStudy (anonymous):

Like I said, you have Sin 30 = apoth/circumradius so circumradius = apoth/Sin 30 = 6/0.5 = 12.

OpenStudy (anonymous):

Oh okay, thanks! Sorry, imma little bit slow today

OpenStudy (anonymous):

np, just makes your life a bit easier if u can reuse something u already know.

OpenStudy (anonymous):

Truth is, all this geometry stuff confuses me.

OpenStudy (anonymous):

A lot of stuff is here http://en.wikipedia.org/wiki/Triangle Comes up all the time, one way or another.

OpenStudy (anonymous):

Thanks you :)

OpenStudy (anonymous):

hi kraken the area A=3(SQRT3)a^2 A=3(SQRT3)(6)^2=187.06 ans